### v1

@@ 3.3 Killing Vectors and Isometries
If a Killing vector $X$ is the infinitesimal generator of a flow $\Psi$, then for two points $p,q\in\mathcal{M}$,
$d(\Psi(t,p),\Psi(t,q))=d(p,q)$
for any value of the flow parameter $t$, so that $\Psi$ is an isometry of the distance measure $d$ induced by the metric tensor $g$. We note that the set of Killing vectors of a metric $g$ form a vector space as well as a Lie algebra (using the Lie bracket for vector fields). In general, Killing vectors of a particular metric $g$ can be calculated, though equation 10 defines a system of overdetermined, linear, homogeneous, partial differential equations known as the Killing equations [11], making the calculation of the Killing vectors of a metric tensor non-trivial. [^2] We note that computational methods have been developed which reduce the complexity of the Killing equations [22]. Herein, such methods are not needed, since the Killing equations for our experiments are sufficiently simple.
Within the context of machine learning, data with continuous-valued features are usually assumed not only to lie in $\mathbb{R}^{n}$, but on the Riemannian manifold $\mathbb{R}^{n}$ equipped with the Euclidean metric tensor, which metric tensor induces the Euclidean distance measure. Within this context, the Killing vectors form a proper subspace of the generators of the affine transformations. However, the Euclidean metric tensor is not always used in machine learning. This occurs when a metric tensor is learned by means of a pullback function induced from a dimensionality reduction algorithm [25]. In section 5.3, we encounter a metric tensor induced by a quadratic feature map, and we restrict our symmetry search space to vector fields which are Killing vectors.
@@ 4 Methods
Our methods assume that a given dataset consists of points residing on a differentiable manifold $\mathcal{M}$, and that a machine learning function $f$ is a smooth function on $\mathcal{M}$. We also assume that the symmetry group we seek is a Lie group generated by 1-parameter transformations known as flows, the infinitesimal generators of which are vector fields which reside in the tangent bundle of $\mathcal{M}$.
There are four components to our methodology. The first component, described in section 4.1, describes the estimation of machine learning functions specific to a dataset. The second component is the estimation of the infinitesimal symmetries of machine learning functions, as well as the estimation of invariant features of a collection of infinitesimal symmetries, and is given in section 4.2. Also in this section are two subsections which speak to the comparison of discovered symmetries to ground-truth symmetries, as well as undesired freedom present in our methods, respectively. The third component of our methodology is isometry discovery, detailed in 4.3. The final component is symmetry enforcement, discussed in section 4.4.
4.1 Estimating Machine Learning Functions
Probability density functions appear in both supervised and unsupervised problems. The output of a logistic regression model or, similarly, a neural network model that uses the softmax function, can be interpreted as estimating the probability that a given point belongs to a particular class. In fact, the probabalistic outputs of many other machine learning models can be analyzed for symmetry, provided that the probability function is differentiable. Invariant features of the symmetries of these functions define level sets upon which the probability is constant. If the flow of an infinitesimal generator can be given, a point in the dataset can be transformed to another point–not generally in the given dataset–and the model should predict the same probability for the transformed point as it would for the original point. Applications include supervised feature engineering and model validation, where predictions on generated points can be used to assess a model’s ability to generalize.
Particularly in the unsupervised case, density estimation can be challenging, complicating methods that require their estimation. A proposed method of addressing the potential shortcoming of density-based symmetry detection is called level set estimation. Level set estimation seeks to learn a function $F:\mathbb{R}^{n}\to\mathbb{R}^{k}$ ($k<n$) such that $F(x_{i})=\mathbf{0}$ for each $x_{i}\in\mathcal{D}$. Any embedded submanifold can be, at least locally, expressed as a level set of a smooth function [15], and so level set estimation operates on the assumption that the data points belonging to a given dataset lie on an embedded submanifold. If such a function can be estimated for a given data set, one can estimate symmetries of the component functions $f_{i}$ of $F$.
Level set estimation can be done as follows, although we note that further development of level set estimation may lead to more effective and/or efficient implementations. First, we construct a function $f$ as a linear combination of pre-determined features, as with a generalized additive model or polynomial regression, with model coefficients $\{a_{i}\}_{i=1}^{m}$. An $m$-component column vector $w$ is constructed using the coefficients $a_{i}$. Then, for each point in the dataset, we set the target to 00, so that a function $f$ is approximated by estimating
(11) ${}Bw=\mathcal{O},$
where $B$ is the feature matrix consisting of $m$ columns and $N$ rows, where $N$ is the number of points in the dataset, and where $\mathcal{O}$ is the zero matrix. OLS regression motivates the way in which we solve Equation (11), though we note that a solution $w_{0}$ is equivalent to a scaling $cw_{0}$ for some constant $c$. We therefore turn to constrained regression, choosing, by default, the constraint $\sum_{i=1}^{m}a_{i}^{2}=1$. Additionally, we note that any loss function can be used, with predictions and targets determined by the left- and right-hand sides of Equation (11), respectively.
We can estimate more than a single component function $f$ by appending additional columns to $w$, in which case we constrain the columns of $w$ to be orthonormal. As appending columns to $w$ is likely to increase any cost function used to estimate $w$, we use an “elbow curve” to determine when the number of columns of $w$ is sufficiently high. That is, in the presence of a sudden comparatively large increase in the cost function as the number of columns of $w$ is increased, the number of columns of $w$ is chosen as the value immediately before the large increase is encountered. We handle the constrained optimization problem of (11) using existing constrained optimization algorithms [27, 18].
The freedom of expression in the components of the level sets is, unfortunately, not fully contained by constrained optimization. When certain functions are chosen as part of the level set estimation, such as polynomials, this scaling issue can extend to multiplication by functions. For example, consider a data set in $\mathbb{R}^{3}$ known a priori to be approximately described by the function $F=\mathbf{0}$, where the components of $f_{1}$ and $f_{2}$ of $F$ are given as
(12) ${}f_{1}=x^{2}+y^{2}-1,\qquad f_{2}=z-1.$
Such a level set could also, at first glance, be described by the following equations:
(13) ${}x^{2}+y^{2}-1=0,\qquad z-1=0,\qquad x(z-1)=0.$
From a mathematical point of view, the third equation is degenerate and does not change the intrinsic dimension of the manifold characterized by this level set, owing to the fact that the rank of this function when $x=0$ is 2. However, within the context of learning level sets from data, the extrinsic dimension may be large, and it may not be practical to examine the components of the learned level set. Furthermore, key points for determining the rank of the corresponding function may not be present in the data.
A potential workaround for this problem is to incrementally increase the allowed degree of the learned polynomial functions in equation 13. For the example considered in equation 12, we can first optimize over the space of degree 1 polynomials, ideally discovering a scaling of the function $f_{2}$. Next, when searching for quadratic components, we can extend the columns of $w$ which correspond to any allowed terms of the form $h(x,y,z)f_{2}$. [^3] At degree 2, $h$ could be, for example, $x$, but not $x^{2}$, since $x^{2}z$ is not a degree 2 polynomial. With constrained optimization, the column space of $w$ can then be expressed as an orthonormal basis. The optimization routine can then continue, with the final level set model for $w$ not including the artificially-created columns. Alternatively, if one is interested primarily in at most quadratic component functions, one can first optimize using affine functions, then project the data onto the lower-dimensional hyperplane defined by the affine components, subsequently searching for quadratic terms on the reduced space where no additional affine components should be present. Experiments which apply these concepts in practice have been given previously [23].
One may well ask whether level set estimation can be accomplished using a neural network rather than constrained polynomial regression. Besides the problem of uniqueness just spoken of, a neural network approach would likely require plentiful off-manifold examples so as not to learn a function which is identically zero. A neural network approach to level set estimation is not explored in this paper, though may be explored in the future.
4.2 Estimating the Infinitesimal Generators and an Associated Invariant Feature Space
@@
With scalar-valued machine learning functions $f_{i}$, we construct a vector-valued function $F$ whose components are defined by the functions $f_{i}$. We can then obtain vector fields which annihilate the components of $F$ by calculating nullspace vectors of $J$, the Jacobian matrix of $F$ [^4] In fact, if $F$ is given in closed form, this can be done using symbolic software and without referencing the original data. More vector fields may annihilate the functions than can be identified by the nullspace of $J$, though our method is inspired by this idea. First, we construct a vector field whose components are an arbitrary linear combination of pre-determined features, as in a general additive model, though we primarily use polynomial features herein. The coefficients of this linear combination reside in a matrix $W$, with the columns of $W$ corresponding to the coefficients for a single vector field. We then estimate $W$ in the following equation:
(14) ${}MW=\mathcal{O},$
@@
where $M=M(\mathcal{J},B)$ is the Extended Feature Matrix computed using the array $\mathcal{J}$ of Jacobian matrices at each point, $B$ is the feature matrix, and $\mathcal{O}$ is the zero matrix. The matrix $M$ is computed via Algorithm 1.
@@
$\mathcal{J}\leftarrow Jacobian(F)(x_{i})$triangleright 3-d array of Jacobian matrices at each point
for i in range($N$do
     $row_{0}\leftarrow padded(B_{i})$triangleright m(n-1) 0’s are appended to the $i^{th}$ row of $B$
     for j in range(1,m-1) do $row_{j}\leftarrow roll(row_{0},j\cdot m)$triangleright Each subsequent row is displaced $m$ entries to the right.
@@
     $mat_{i}\leftarrow\mathcal{J}_{i}b_{i}$triangleright Multiply the Jacobian of $F$ at $x_{i}$ by the matrix $b_{i}$
@@
$M\leftarrow StackVertical([mat_{0},\dots mat_{N-1}])$triangleright This matrix has size $nN\times nm$.
@@
As with Equation (11), a solution to Equation (14) is estimated using constrained optimization of a selected loss function, for which we turn to McTorch [18]. Using this pytorch-compatible library, one can choose from a variety of supported loss functions and (manifold) optimization algorithms. As with level set estimation, we sequentially increase the number of columns of $W$ until an “elbow” in cost function values is reached.
@@
However, we find the same unwanted freedom as with level set estimation, since $X(f)=0$ implies that for a smooth function $h$, $hX(f)=0$. In the generic case, it is commonly assumed that the components of $X$ are polynomial functions, in which case it is possible that a such a pair $X$ and $hX$ may exist in the search space. The means of dealing with this unwanted freedom can depend on the particular experiment: if there are no affine symmetries, for example, vector fields with components of degree $2$ can be approximated without regard to this issue: this is further expounded upon in section 4.2.2.
To define a model which is invariant with respect to the flows of pre-computed vector fields $X_{i}$, we seek features $h^{j}$ (separate of the functions $f_{i}$) such that $X_{i}(h^{j})=0$. This we do by estimating $v$ in the following equation:
(15) ${}M_{2}v=\mathcal{O},$
where $M_{2}=M_{2}(\mathcal{J},B,\mathcal{B})$ is the Invariant Function Extended Feature Matrix computed via Algorithm 2, $\mathcal{B}$ is the feature matrix of $W$ (already estimated), and $B$ is the feature matrix of the unknown function.
Algorithm 2 Constructing the Invariant Function Extended Feature Matrix $M_{2}$
$m\leftarrow$ number of features
$n\leftarrow$ dimension of space
$N\leftarrow$ number of points in the dataset
$\bar{J}(B)(x_{i})\leftarrow$ Jacobian matrix of features of $B$ at $x_{i}$triangleright size $m\times n$: derivatives of raw features.
$\mathcal{J}\leftarrow$ $[\bar{J}(B)(x_{i})]$triangleright Array of Feature Jacobian matrices at each point.
$M=M(\mathcal{J},\mathcal{B})$triangleright Extended Feature matrix computed via Algorithm 1
$M_{2}=MW$$\triangleright$ This is not $\mathcal{O}$, since the arguments of $M$ are different than in 14
This method assumes that the invariant functions can be expressed in terms of prescribed features, which, in our case, are typically polynomial functions. The quantity $v$ can be estimated from Equation 15 in much the same way that we estimate other quantities, in that we constrain $v$ to be orthogonal, then optimize a selected loss function using McTorch [18]. We should note that while this method may produce invariant features, they are not guaranteed to be functionally independent. For example, in our polynomial feature space, the vectors corresponding to $x$ and $x^{2}$ are orthogonal, though the functions themselves are not functionally independent.
4.2.1 Comparison of Discovered Symmetries to Ground-Truth Symmetries
To evaluate symmetry discovery methods in general, it is important to quantify the ability of a given method to recover ground truth symmetries in controlled experiments. For a ground truth vector field $X$ and an estimated vector field $\hat{X}=\sum_{i=1}^{N}\hat{f}_{i}\partial_{x^{i}}$, we define the angle between them by
[…]
where $\langle X,\hat{X}\rangle_{g}=\sum_{i,j}f_{i}\hat{f}_{j}g_{ij}$, $||X||_{g}=\sqrt{\langle X,X\rangle_{g}}$, and where
[…]
with the region $\Omega$ being defined by the range of a given dataset–ordinarily, this is taken to be the full range of the dataset.
@@ 4.2.1 Comparison of Discovered Symmetries to Ground-Truth Symmetries
4.2.2 A Note about the Freedom Present in Identifying Vector Field Symmetries
@@
For example, a circle centered at the origin, characterized by $F=0$ where $F(x,y)=x^{2}+y^{2}-r^{2}$ for some real number $r$, exhibits rotational symmetry described by $X=-y\partial_{x}+x\partial_{y}$. However, $\frac{1}{y}X(F)=0$, so that the flow of the vector field $-\partial_{x}+\frac{x}{y}\partial_{y}$, where defined, is also a symmetry of the circle.
@@
This may seem to introduce a theoretical problem requiring greater care when detecting non-affine symmetry. However, our proposed methods of constructing/training models which are invariant with respect to the symmetry group requires only the identification/learning of functions which are invariant with respect to the vector fields. Thus, both $X$ and $fX$ are valid answers for “ground truth” symmetries, since a function $h$ is $X$-invariant if and only if it is $fX$-invariant.
We note that this lack of uniqueness does present a challenge when estimating suitable vector fields. When estimating the symmetries via constrained regression, $X$ and $fX$ may or may not both be present in the search space. As mentioned in section 4, one potential workaround is owing to the fact that symmetry estimation can be done symbolically. Consider the example of $F=0$ given previously in this section. The Jacobian matrix of $F$ is given as
$J=\begin{bmatrix}2x&2y&2z\end{bmatrix},$
and a basis for the nullspace of $J$ (using functions as scalars) is
$\left\{\begin{bmatrix}-y\\ x\\ 0\end{bmatrix},\begin{bmatrix}0\\ -z\\ y\end{bmatrix}\right\},$
@@
which vectors correspond to tangent vectors $-y\partial_{x}+x\partial_{y}$ and $-z\partial_{y}+y\partial_{z}$, respectively. The flows of these vector fields correspond with rotations (about the origin) in the $(x,y)$ and $(y,z)$ planes, respectively. However, only two vector fields can be recovered using this method, which in this case has neglected another symmetry, namely $-z\partial_{x}+x\partial_{z}$, corresponding to rotations in the $(x,z)$ plane. In fact, each of the three vector fields corresponding to rotations about a coordinate axis are isometries of the sphere described by $F=0$. We should note, however, that the generator for rotations in the $(x,z)$ plane can be expressed as a linear combination of the other rotations using functions as scalars:
$\dfrac{z}{y}\left(-y\partial_{x}+x\partial_{y}\right)+\dfrac{x}{y}\left(-z\partial_{y}+y\partial_{z}\right)=-z\partial_{x}+x\partial_{z}.$
Therefore, it is plausible that applications exist in which a symbolic approach to continuous symmetry discovery may be sufficient. Our approach does not rely on symbolic software, however.
@@
We also mention that the issue of unwanted freedom can be addressed by reducing the search space of vector fields to linear combinations of geometrically significant vector fields. In section 4.3, we discuss the restriction of the search space to Killing vectors.
@@
Another point which may help to address concerns about our handling of the difficult nature of non-affine symmetry detection is in the relationship between the flows of $X$ and $fX$ generally. The trace of the flow of $X$ through the point $p$ is characterized by the level set $h_{i}=c_{i}$, where $\{h_{i}\}$ is a complete set of $X$-invariant functions. Since a function is $X$-invariant if and only if it is $fX$-invariant, a complete set of invariant functions for $fX$ can be taken to be $\{h_{i}\}$ without loss of generality. Thus, the level set $h_{i}=c_{i}$ also characterizes the trace of the flow of $fX$ through the point $p$. This argument assumes that the flows of $X$ and $fX$, as well as the vector fields themselves, are well-defined in an open neighborhood about the point $p$.
In fact, in our example above with $F=x^{2}+y^{2}-r^{2}$, the flow of $-\partial_{x}+\frac{x}{y}\partial_{y}$, assuming $y>0$, is given as
$\Phi(t,(x,y))=\left(x-t,\sqrt{y^{2}+2tx-t^{2}}\right),$
@@
which trace is (part of) a circle, and we note that the flow parameter is $-x$. The trace of this flow through a point is equivalent to the trace of $X$ through the same point.
4.3 Isometry Discovery
As mentioned in section 4.2.2, the fact that a function is $X$-invariant if and only if it is also $fX$-invariant can lead to complications when estimating vector field symmetries. One way of attempting to handle this issue is to restrict the search space of symmetries so that, generally, such multiples are not present in the search space. In this section, we outline a technique to restrict the search space to isometries of a distance measure induced by a Riemannian (or pseudo-Riemannian) metric tensor.
Our model assumption is that our data lies on a manifold $\mathcal{M}$ equipped with a metric tensor $g$, and that a machine learning function $f$ on $\mathcal{M}$ is smooth. As stated before, a basis for the vector space of Killing vectors $\{X_{i}\}_{i=1}^{s}$ can be computed by solving equation 10, which is a system of overdetermined, first order, homogeneous partial differential equations known as the Killing equations. Once a basis is identified, we define $X$ to be an arbitrary linear combination of the basis vector fields:
$X=\sum_{i=1}^{s}c_{i}X_{i},$
then optimize the real coefficients $\{c_{i}\}_{i=1}^{s}$, subject to $\sum_{i=1}^{s}c_{i}^{2}=1$, in the equation $X(f)=0$. This is a case of equation 14 in which the feature matrix $B_{sr}$ is defined by the components of the Killing vectors, $\mathcal{J}$ is the Jacobian of $f$, and the single column of $W$ is defined by the coefficients $c_{i}$. To search for multiple Killing vectors which annihilate $f$, additional columns can be appended to $W$.
4.4 Symmetry Enforcement
We have discussed how, for a given smooth machine learning function $f$, one may estimate vector fields $X_{i}$ for which $X_{i}(f)=0$ identically. We now turn to the topic of enforcing continuous symmetries in a smooth machine learning model during training using vector fields.
Suppose that, in the usual way, one seeks to learn a function $f$ mapping data instances $x_{i}$ to targets $y_{i}$, where $x_{i}\in\mathbb{R}^{d}$ and $y_{i}\in\mathbb{R}^{m}$. Suppose also that the function $f$ is estimated by means of the minimization of $L(f(\mathbf{x}),\mathbf{y})$ for a smooth loss function $L$.
Now suppose that it is desired that the function $f$ learns to exhibit continuous symmetry with respect to the independent infinitesimal generators $\{X_{k}\}_{k=1}^{s}$: that is, for each component $f_{j}$ of $f$,
$X_{k}(f_{j})=0$
for $1\leq k\leq s$ and $1\leq j\leq m$. Our method of symmetry enforcement in training is to minimize the following loss function:
The loss function given in 17 bears striking resemblance to that of a PINN, where our $\tilde{L}$ term corresponds to the appendage of differential constraints imposed by the governing equations for a given physical system [20]. Adding to this similarity is the fact that our symmetry enforcement term is a differential constraint, so that our method of symmetry enforcement appears to be a special case of a PINN in which the differential operators are linear, first-order, and homogeneous.
Despite the mathematical equivalence of enforcing continuous symmetries using vector fields and the enforcement of differential equations governing a physical system, we note that PINNs do not erase the need for symmetry enforcement using vector fields. For example, consider the enforcement of rotational symmetry in an image classification task: a symmetry-based approach provides a relatively clear path forward, since a vector field corresponding to planar rotational symmetry can readily be given. Such a task is apparently more cumbersome using a PINN, where underlying differential constraints are derived from the description of desired symmetry. Thus, a symmetry-based regularization is more appropriate when the desired symmetries can readily be given or, perhaps, more easily estimated than the full set of differential equations underlying a physical system. On the other hand, if the differential equations underlying a given system are fully known, regularization using a PINN-based approach may be the most appropriate method.

### v2

@@ 3.4 Killing Vectors and Isometries
If a Killing vector $X$ is the infinitesimal generator of a flow $\Psi$, then for two points $p,q\in\mathcal{M}$, $d(\Psi(t,p),\Psi(t,q))=d(p,q)$ for any value of the flow parameter $t$, so that $\Psi$ is an isometry of the distance measure $d$ induced by the metric tensor $g$. We note that the set of Killing vectors of a metric $g$ form a vector space as well as a Lie algebra (using the Lie bracket for vector fields). Killing vectors of a particular metric $g$ can be calculated by solving Equation 9, which defines a system of overdetermined, linear, homogeneous, partial differential equations known as the Killing equations [Gover1], in general numerically.  [^2] Computational methods have been developed which reduce the complexity of the Killing equations [Shaw2]. Herein, such methods are not needed, since the Killing equations for our experiments are sufficiently simple.
In the context of machine learning, data with continuous-valued features are usually assumed not only to lie in $\mathbb{R}^{n}$, but on the Riemannian manifold $\mathbb{R}^{n}$ equipped with the Euclidean metric tensor, which induces the Euclidean distance measure. In this context, the Killing vectors form a proper subspace of the generators of the affine transformations. However, the Euclidean metric tensor is not the only one used in machine learning: e.g., a different metric tensor may be learned by means of a pullback function induced from a dimensionality reduction algorithm [Sun1]. In such contexts, the Killing vectors may generate non-affine transformations. In Section 6.3, we encounter a metric tensor induced by a quadratic feature map, and we restrict our symmetry search space to vector fields which are Killing vectors.
4 Problem Formulation
Here we precisely formulate the symmetry discovery problem in a way that will inspire our proposed solution. A data-generating distribution $\mathcal{D}$ on a differentiable manifold $\mathcal{M}$ embedded in $\mathbb{R}^{n}$ is fixed but unknown to us. A model class $\mathcal{H}$ consisting of vector fields is also fixed. For instance, $\mathcal{H}$ may be the set of vector fields with coefficient functions that are polynomials with degree at most $d$. For $X\in\mathcal{H}$ and a smooth function $f:\mathbb{R}^{n}\to\mathbb{R}$, we denote by $\ell_{f}(X,z)=\ell(X[f](z))$ the loss of $X$ at the point $z$ (e.g., $\ell_{f}(X,z):=X[f](z)^{2}$). We then define the risk of $X$ as
(10) $R_{f}(X):=\mathbb{E}_{z\sim\mathcal{D}}[\ell_{f}(X,z)].$
Given a dataset $D:=\{x_{1},...,x_{r}\}$ sampled independently from $\mathcal{D}$ and a machine learning function $f$, our goal is to identify infinitesimal generators $X\in\mathcal{H}$ that minimize $R_{f}(X)$.
Formally, we define a symmetry learner $L$ to be a function from datasets to elements of $\mathcal{H}$. We say that $L$ is $(\epsilon,\delta)$-probably approximately correct (PAC) if there is some $r_{0}(\epsilon,\delta)$ such that, given $r\geq r_{0}(\epsilon,\delta)$ data samples, for all data-generating distributions $\mathcal{D}$, with probability at least $1-\delta$ over sampled datasets, $L$ outputs an $X$ with risk satisfying the following:
[…]
Our two goals, then, are symmetry discovery and symmetry enforcement.
While the estimation of machine learning functions is required prior to symmetry discovery, the novelty contained in the conference paper for which this is an extension dealt primarily with level set estimation, which is not the subject of the experiments contained herein. We therefore refer the reader to the original conference paper [Shaw1] for information regarding the estimation of machine learning functions.
@@ 5 Methods
There are several components to our methodology. First is the estimation of the infinitesimal symmetries of machine learning functions, detailed in Section 5.1. Another component of our methodology is isometry discovery, detailed in Section 5.2. In Section 5.3, we remark on the ill-posedness of naïve formulations of the symmetry discovery problem. In Section 5.4, we give our results analyzing the robustness of symmetry discovery to error in estimation of the machine learning function. We describe how we evaluate our methods by comparing discovered symmetries to ground-truth symmetries in Section 5.5. In Section 5.6, we present two techniques for enforcement of discovered symmetries in downstream machine learning models.
5.1 Estimating the Infinitesimal Generators
@@
We first describe the intuition behind our method. With scalar-valued machine learning functions $f_{i}$, we construct a vector-valued function $F$ whose components are defined by the functions $f_{i}$. We can then obtain vector fields which annihilate the components of $F$ by calculating nullspace vectors of $J$, the Jacobian matrix of $F$. More vector fields may annihilate the functions than can be identified by the nullspace of $J$, since some vector fields, though linearly independent over the field of real numbers, are not linearly independent over the ring of $\mathcal{C}^{\infty}$ functions. However, our method is inspired by the notion of looking to the nullspace of $J$. First, we construct a model class of vector fields whose components are norm-constrained linear combinations of pre-determined features, as in a general additive model, though we primarily use polynomial features herein. The advantage of the choice of polynomial features is that this results in a nested sequence of model classes (indexed by the maximum polynomial feature degree) that is dense in the set of smooth vector fields. The coefficients of the linear combination of a given model reside in a matrix $W$, with the columns of $W$ corresponding to the coefficients for a single vector field. We then estimate $W$ by solving the following minimization problem, which is constrained by the coefficient norm bound defining the model class:
(12) $W:=\operatorname*{arg\,min}_{W_{*}}\ell(MW_{*}),$
@@
where $M=M(\mathcal{J},B)$ is the Extended Feature Matrix computed using the array $\mathcal{J}$ of Jacobian matrices at each point, $B$ is the feature matrix, and $\ell(\cdot)$ is the sum of losses over all columns of $W_{*}$. The matrix $M$ is computed via Algorithm 1.
@@
$\mathcal{J}\leftarrow Jacobian(F)(x_{i})$ $\triangleright$ 3-d array of Jacobian matrices at each point
for i in range($N$) do
  $row_{0}\leftarrow padded(B_{i})$ $\triangleright$ m(n-1) 0’s are appended to the $i^{th}$ row of $B$
  for j in range(1,m-1) do $row_{j}\leftarrow roll(row_{0},j\cdot m)$ $\triangleright$ Each subsequent row is displaced $m$ entries to the right.
@@
  $mat_{i}\leftarrow\mathcal{J}_{i}b_{i}$ $\triangleright$ Multiply the Jacobian of $F$ at $x_{i}$ by the matrix $b_{i}$
@@
$M\leftarrow StackVertical([mat_{0},\dots mat_{N-1}])$ $\triangleright$ This matrix has size $nN\times nm$.
@@
A solution to Equation (12) is estimated using constrained optimization of the selected loss function, for which we turn to McTorch [mctorch]. Using this PyTorch-compatible library, we can choose from a variety of supported loss functions and (manifold) optimization algorithms. To decide on the number of columns of $W$, we sequentially increase this number until an “elbow” in the cost function values is reached. Despite claims to the contrary [hu1], this approach is amenable to automatic discovery of the number of symmetries, due to existing methods of automatic “elbow point” discovery [Onumanyi1].
@@
However, we find the same unwanted freedom as with level set estimation [Shaw1], since $X(f)=0$ implies that for any smooth function $h$, $hX(f)=0$. We commonly assume that the components of $X$ are polynomial functions, in which case it is possible that a such a pair $X$ and $hX$ may exist in the search space. There are a few ways of dealing with this issue: primarily, by reducing the search space (to, say, Killing vectors) so that $X$ and $hX$ are not generally both in the search space. This issue is further expounded upon in Section 5.3.
5.2 Isometry Discovery
As mentioned in Section 5.3, the fact that a function is $X$-invariant if and only if it is also $fX$-invariant can lead to complications when estimating vector field symmetries. One way to handle this issue is to restrict the search space of symmetries so that such multiples are not present in the search space. In this section, we outline a technique to restrict the search space to isometries of a distance measure induced by a Riemannian (or pseudo-Riemannian) metric tensor.
Our model assumption is that our data lies on a manifold $\mathcal{M}$ equipped with a metric tensor $g$, and that a machine learning function $f$ on $\mathcal{M}$ is smooth. As stated before, a basis for the vector space of Killing vectors $\{X_{i}\}_{i=1}^{s}$ can be computed by solving Equation 9, which is a system of overdetermined, first order, homogeneous partial differential equations known as the Killing equations. Once a basis is identified, we define $X$ to be an arbitrary linear combination of the basis vector fields: $X=\sum_{i=1}^{s}c_{i}X_{i}.$ We then optimize the real coefficients $\{c_{i}\}_{i=1}^{s}$ in the equation $X(f)=0$ subject to the constraint $\sum_{i=1}^{s}c_{i}^{2}=1$. This is a case of Equation 12 where the feature matrix $B_{sr}$ is defined by the components of the Killing vectors, $\mathcal{J}$ is the Jacobian of $f$, and the single column of $W$ is defined by the coefficients $c_{i}$. To search for multiple Killing vectors which annihilate $f$, one appends additional columns to $W$.
@@ 5.2 Isometry Discovery
5.3 A Note about the Freedom Present in Identifying Vector Field Symmetries
@@
For example, a circle centered at the origin, characterized by $F=0$ where $F(x,y)=x^{2}+y^{2}-r^{2}$ for some real number $r$, exhibits rotational symmetry described by $X=-y\partial_{x}+x\partial_{y}$. However, $\frac{1}{y}X(F)=0$, so that the flow of the vector field $\partial_{x}+\frac{x}{y}\partial_{y}$, where defined, is also a symmetry of the circle.
@@
This may seem to introduce a theoretical problem requiring greater care when detecting non-affine symmetry. However, our proposed methods of constructing/training models which are invariant with respect to group actions requires only the identification/learning of functions which are invariant with respect to the vector field infinitesimal generators. Thus, both $X$ and $fX$ are valid answers for “ground truth” symmetries, since a function $h$ is $X$-invariant if and only if it is $fX$-invariant.
@@
This non-uniqueness presents a challenge when estimating suitable vector fields. When estimating the symmetries via constrained regression, $X$ and $fX$ may or may not both be present in the search space. As mentioned in Section 5, one potential workaround is to do symmetry estimation symbolically. Consider the example of $F=0$ given previously in this section. The Jacobian matrix of $F$ is given as $J=\begin{bmatrix}2x&2y&2z\end{bmatrix}$, and a basis for the nullspace of $J$ (using functions as scalars) is $\{[-y,x,0]^{T},[0,-z,y]^{T}\}$, which vectors correspond to tangent vectors $y\partial_{x}+x\partial_{y}$ and $z\partial_{y}+y\partial_{z}$, respectively. The flows of these vector fields correspond with rotations (about the origin) in the $(x,y)$ and $(y,z)$ planes, respectively. However, only two vector fields can be recovered using this method, which in this case has neglected another symmetry, namely $z\partial_{x}+x\partial_{z}$, corresponding to rotations in the $(x,z)$ plane. In fact, each of the three vector fields corresponding to rotations about a coordinate axis are isometries of the sphere described by $F=0$. We should note, however, that the generator for rotations in the $(x,z)$ plane can be expressed as a linear combination of the other rotations using functions as scalars: $\dfrac{z}{y}\left(-y\partial_{x}+x\partial_{y}\right)+\dfrac{x}{y}\left(-z\partial_{y}+y\partial_{z}\right)=-z\partial_{x}+x\partial_{z}.$ Thus, it is plausible that applications exist in which a symbolic approach to continuous symmetry discovery may be sufficient. Our approach does not rely on symbolic software, however.
@@
The issue of unwanted freedom can be addressed by reducing the search space of vector fields to linear combinations of geometrically significant vector fields. In Section 5.2, we discuss the restriction of the search space to Killing vectors.
@@
Another point which may help to address concerns about our handling of the difficult nature of non-affine symmetry detection is in the relationship between the flows of $X$ and $fX$ generally. The trace of the flow of $X$ through the point $p$ is characterized by the level set $h_{i}=c_{i}$, where $\{h_{i}\}$ is a complete set of $X$-invariant functions. Since a function is $X$-invariant if and only if it is $fX$-invariant, a complete set of invariant functions for $fX$ can be taken to be $\{h_{i}\}$ without loss of generality. Thus, the level set $h_{i}=c_{i}$ also characterizes the trace of the flow of $fX$ through the point $p$. This argument assumes that the flows of $X$ and $fX$, as well as the vector fields themselves, are well-defined in an open neighborhood of the point $p$.
@@
In fact, in our example above with $F=x^{2}+y^{2}-r^{2}$, the flow of $\partial_{x}+\frac{x}{y}\partial_{y}$, assuming $y>0$, is given as $\Phi(t,(x,y))=\left(x-t,\sqrt{y^{2}+2tx-t^{2}}\right),$ which trace is (part of) a circle with flow parameter x-x. The trace of this flow through a point is equivalent to the trace of $X$ through the same point.